Given the following data, what is the Mean Absolute Deviation (MAD):

Month Demand Forecast
Jan 100 110
Feb 100 100
Mar 120 100
Apr 110 90
May 100 110
Jun 90 100
Jul 80 90
Aug 90 80
Sep 100 110
Oct 110 100
Nov 110 110
Dec 120 110
A.

10

B.

20

C.

30

D.

133.33
Answer #1

calculation of MAD: Month Demand forcast Deviation A-f Absolute deviation JA-EL -10 H-10=10 1201=20 | 120l = 20 1-10laco 11-1MAD- MAD= {{A-FI / A: Demand A= Demand F = Pore Cast 120 MADELO MADE 10 12 | MAD=10) option ® comect.